1. Which of the following BEST describes a company's proper liquidity management? Multiple Choice Liquitity management...

1.

Which of the following BEST describes a company's proper liquidity management?

Multiple Choice

  • Liquitity management is a balancing act; managers try to find liquidity levels that are neither too high not too low.

  • A Financial Manager will try to keep as much cash on the books as possible to maximize short-term earnings.

  • A company should never keep cash in its account because bond coupon payments can be deferred for up to a year without penalty.

  • Liquidity levels that are too low will never cause a firm to go bankrupt.

2

When the corporate tax structure features multiple tax brackets each with different tax rates, which of the following should a financial manager use in deciding whether to undertake a corporate spending project?

Multiple Choice

  • Average tax rate

  • Marginal tax rate

  • When there are multiple tax brackets with different tax rates, average and marginal tax rates will always be the same.

3.

Jim's Food Mart Inc.'s balance sheet in 2017 showed Current Assets of $15,000 and Current Liabilities of $9,600. In 2018 these figures were $15,675 and $11,290, respectively. What was the company's change in Net Working Capital in 2018?

Multiple Choice

  • +$4,385

  • +$1,015

  • +$5,400

  • -$1,015

  • -$985
Answer #1

1) If the liquidity is kept at high level under the fear of not being capable of meeting financial requirements in time and funds available are not invested in sure to count on losses for no returns on funds available.

* If funds invested without care for even minimum requirement of cash/ liquidity, in case of urgent need the financial commitments May may not meet deadline and result in losses in the form of penalty/ high rate of interest.

Answer:A

Liquidity management is balancing act,managers try to find liquidity levels that are neither too high not too low.

2) Average tax measure tax burden while marginal tax rates measure impact of taxes on incentives to earn,save ,invest or spend.

Answer: A

When there are multiple tax brackets with multiple tax rates,we use average tax rate.

3) Change in net working capital= working capital (2018)- working capital (2017)

Working capital (2018)= current assets - current liabilities

= 15675-11290= 4385

Working capital (2017) = 15000- 9600= 5400

Net change in working capital = 4385- 5400 = -1015